OpenStudy (anonymous):

Show that p=q(x+y/xy) is a sol'n to the partial diff. equation x^2 dp/dx - y^2 dp/dy = 0. This isn't true... right?

OpenStudy (anonymous):

(x+y)/(xy)...?

OpenStudy (anonymous):

when i calculated them, i got 2x^3y-2xy^3 = 0

OpenStudy (anonymous):

= x/xy + y/xy = 1/y + 1/x q*d/dx { ((1/y+1/x) }= -q/x^2 q*d/dy { ((1/y+1/x) }= -q/y^2 x^2 (-q/x^2) - y^2 (-q/y^2) = ?

OpenStudy (anonymous):

hmmm ok.